a lamp with a circular base 14cm in diameter was on a square mat with an area of 196 cm^2. How many cm^2 of the mat were uncovered?
Answer by oberobic(2304)

A lamp with a circular base of 14 cm will have a radius of 7 cm.
The area will be pi*r^2 = 49 pi = 153.9
.
The area of the square mat is 196 cm^2..
Each side = sqrt(196) = 14 cm.
.
196-153.9 = 42.1 cm^2 will be uncovered.
.